HUBBLE SPACE TELESCOPE DAILY REPORT #828 PERIOD COVERED: 0800 EST Daily Status Report as of 053/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 3372 (GHRS Point Spread Function Determination) The High Resolution Spectrograph was used to make four observations of standard star HD141556 to help measure the point spread function of the large science aperture. All four observations were received. The first and the third observations were deflection calibrations. The other two were the spiral search and the peakup. The target was found in the ninth map. After the peakup the target flux was 45044. 1.2 Completed FOS/RD/BL 4115 (Imaging and Spectroscopy of Nearby Quasars) The Faint Object Spectrograph was used to make three spectral observations on both the red and blue side of optically bright quasar PG1116+215 whose visual magnitude is around 15. The observations were received and displayed. They were reported to look normal. 1.3 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to make two high latitude observations taken in parallel with the prime science observations. The observations were received and displayed in the OSS. The observations were reported to look okay. 1.4 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Two observations were made using both the Wide Field Camera and the Planetary Camera to make internal flat calibration images. This is to check for new forms of contamination (i.e., Measles and Daisies), and the return of quantum efficiency hystoresis. Both observations were executed as scheduled, and no problems were reported. 1.5 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make observations of the protoplanetary system PK147-02D1. The observations were executed as scheduled, and no problems were reported. 1.6 Completed GHRS/2 3626 (Empirical Determination of Wind Velocity and Density Laws) The High Resolution Spectrograph was used to make spectral observations of the eclipsing binary star Zeta Auriega (aka, HD32068) during part of its eclipse phase. The observations were executed as scheduled, and no problems were reported. 1.7 Completed GHRS/2 3706 (Nature of Gas Loops in Milky Way Halo) The High Resolution Spectrograph was used to make observations of high galactic latitude star HD+49D2137 to study the neutral hydrogen clouds which are moving toward the plane of our galaxy. The observations were monitored in the OSS, and no problems were reported. 1.8 Completed GHRS/2 4067 (GHRS 1st Order Grating Spectral Lamp Test) The High Resolution Spectrograph was used to make spectral calibration lamp mini-function tests using the first order gratings. The calibration observations were executed as planned, and monitored in the OSS. No problems were reported. 1.9 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.10 Completed GHRS/2 4768 (Dynamics and Energy Balance of Interstellar Transition Regions The High Resolution spectrograph was used to make spectral observations of the star Beta Draconis (aka, HD159181) to look for the redshift of spectral lines of carbon, oxygen, and Silicon. The observations were monitored by the OSS, and no problems were reported. 1.11 Completed PC 4347 (WF/PC calibration: Flood Retention Test) The Planetary Camera was used to make Earth Calibration images using filter F439W to determine how much of the last ultraviolet flood still remains in the system. The observations were monitored in the OSS, and no problems were reported. 1.12 Completed GHRS/2 4792 (Hydrogen Lyman Alpha Dayglow Emission Line Profiles) The High Resolution Spectrograph was used to make two line profiles of the Lyman Alpha dayglow of the central meridian of the planet Jupiter. This was at the equator and at a mid latitude. Both observations were received and displayed. These were a wave length calibration and spectral observations. They were reported to look okay. 1.13 Completed PC 4712 (WF/PC Calibration: Measles Monitor) The Planetary Camera was used to make three Earth flat calibration images using filter F517N. No problems were seen with the telemetry. 1.14 Completed GHRS/2 4766 (Boron in Population II Dwarfs, Part 2) The High Resolution Spectrograph was used to make a large series of observations (24) of the metal poor unresolved dwarf star HD140283. The observations were received, and where possible displayed. Most were reported to look normal. One of the observations was affected by spacecraft jitter during a night/day transition. 1.15 Completed GHRS/2 3463 (Highly Ionized Nitrogen in Galactic Halo) The High Resolution Spectrograph was used to make nine observations of the active galactic nuclei of NGC-3783. All nine observations were received and displayed. The confirmation image showed the target well centered. One of the spectral observations showed several observer bias disturbances. The spectral observations were reported to look okay. 1.16 Completed FOC/96 (3591 (Mass Stars in Starburst Galaxies- FOC Images (Part 1)) The Faint Object Camera (f/48) was used to make one ultraviolet observation of the star burst formation galaxy NGC-5253. The observations was received and displayed. It was reported to look okay. 1.17 Completed WFC 4713 (Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make one internal flat observations looking of quantum efficiency changes which may have occurred during decontamination operations. The observations was received and displayed. It was reported to look okay. 1.18 Completed PC 3769 (Structure of Inner Coma of Comet Chiron) The Planetary Camera was used to make seven observations with filter F555W of the object/comet known as 2060 Chrion. This was to study its the outer edge (i.e., the exopause) of the gravitationally bound dust cloud around its nucleus (i.e., the inner coma). All seven observations were received, but only the first four were displayed. They looked okay. No other problems were reported. 1.19 Completed WFC 4713 (WF/PC Calibration: Internal Delta Flats) One internal flat calibration observations was made using the Wide Field Camera. It was received and displayed. It was reported to look normal. 1.20 Initiated WFC/PC 4207 (WF/PC Calibration: Photometric Calibration Monitor) Both the Wide Field Camera and the Planetary Camera were used to make measure the ultraviolet flux of standard star BD+75D325 to measure the unit's throughput. The observations were monitored in the OSS, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 34 Successful : 34 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Significant HSTARs: None 2.4 Operations Note: The special FGS #2 diagnostic test was conducted last Friday (Feb. 19, 1993) The actions of the FGS were closely monitored while Star Selector Servo A was rotated ten turns in each direction at a rate of 10,240 arcsecond per second, to help redistribute the lubricant in the unit. Because the torque values were reaching saturation near the end of the test, the test was terminated until further data analysis can be performed. The star selectors were then moved to a safe location, and powered down. A detailed analysis of the anomaly is currently under way. A special review meeting has been called for Feb. 23, 1993, to review all data. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC/PC 4207 (Photometeric Calibration Monitor) 3.2 Complete PC 4714 (WF/PC Calibration: Internal Delta Flats) 3.3 Complete Two Set of FGS 4064 (Crit. Test of Galactic Escape Velocity at R(Sun)) 3.4 Complete Three Sets of WFC 4713 (WF/PC Calibration: Internal Delta Flats) 3.5 Complete WFC 4105 (HST Medium Deep Survey: High Latitude) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete GHRS/2 4068 (GHRS Detector Calibration) 3.8 Complete PC 3769 (Structure of the Inner Coma of Comet Chiron) 3.9 Initiate FOS/BL 3545 (Evolution of the UV Spectra in Early Type Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-